Calculate the volume of N2 that results when 0.109 L of the gas is heated from 297.0 K to 332.0 K at 3.00 atm


1
Expert's answer
2021-03-12T06:01:08-0500

Since the pressure is constant, the combined gas law is:

V1/T1 = V2/T2

where V1 - initial volume, T1 - initial temperature, V2 - final volume, T2 - final temperature.

From here:

V2 = V1 × T2 / T1 = 0.109 L × 332.0 K / 297.0 K = 0.122 L


Answer: 0.122 L